OpenStudy (anonymous):

Find all points of intersection: y=-x^2+4x and y=x^2

OpenStudy (anonymous):

Equate both the equations to each other. -x^2+4x=x^2 -2x^2+4x=0 2x(4-x)=0 x=0,4. Point of y is for x =0 is y = 0

OpenStudy (precal):

Sorry you factored it incorrectly. 2x(-x+2)=0 x=0 and x=2 (0,0) and (2,4) are the points of intersection.